# Perfect Quantum State Revivals: Designing Arbitrary Potentials with Specified Energy Levels

### v1

@@ Abstract
It is known that there exist a limited number of analytic potentials with the unusual property that any bound quantum state therein will be periodic in time. This is known as a perfect quantum state revival. Examples of such potentials are the infinte well, quantum harmonic oscillator and the Pöschl-Teller potentials; here, we present a general method of designing such potentials. A key requirement is that their energy eigenvalues have integer spacings (up to a prefactor). We first analyze the required conditions which permit quantum state revivals for potentials in general, and then we use techniques of iterated Hamiltonian intertwining to construct potentials exhibiting perfect quantum revivals. Our method can readily be extended to multiple dimensions.
@@ III Constructing potentials with desired spectra
The technique begins with a known potential $V_{0}(x)$ and associated Hamiltonian,
@@
A new potential $V_{1}(x)$ shall be constructed such that it has all the energy levels that $V_{0}(x)$ has, plus an additional level that shall lie below all those of $V_{0}(x)$. The total number of bound states of $V_{1}(x)$ is thus one greater than in $V_{0}(x)$. If $V_{0}(x)$ is one of the potentials whose discrete levels obey the contraints of (2) and we then add a new energy level below its ground state at the right energy, a new potential retaining the quantum state revival property will result. It is then possible to iterate the intertwining procedure with $V_{1}(x)$ as the starting point to add another energy level below its ground state, and then continue to create as many new energy levels as desired. In the standard intertwining procedure as described in [10, 11, 8, 12], an operator $A^{+}_{1}$ is sought which connects two Hamiltonians $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$ such that
$A^{+}_{1}\mathcal{H}_{1}=\mathcal{H}_{0}A^{+}_{1}\,$. (4)
In the first-order intertwining, which we shall limit ourselves to, the operator $A_{1}^{+}$ has the following form:
$A^{\pm}_{1}=\frac{1}{\sqrt{2}}\left(\pm\frac{d}{dx}+U_{1}(x)\right)\,$ (5)
The function $U_{1}(x)$ is to be determined, and $A^{-}_{1}=(A^{+})^{\dagger}$. We take note that $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$ can be factored, where $\mathscr{E}$ is an undetermined constant:
$\mathcal{H}_{0}=A^{+}_{1}A^{-}_{1}+\mathscr{E}\,,\qquad\mathcal{H}_{1}$ $=A^{-}_{1}A^{+}_{1}+\mathscr{E}\,$. (6)
If we now take $E_{1}$ as an eigenenergy of $\mathcal{H}_{1}$ and $\psi_{1}$ as its normalized eigenfunction, then $A^{-}_{1}A^{+}_{1}\psi_{1}=(E_{1}-\mathscr{E})\psi_{1}$. Multiplying from the left by $A^{+}_{1}$ results in $A^{+}_{1}A^{-}_{1}(A^{+}_{1}\psi_{1})=(E_{1}-\mathscr{E})(A^{+}_{1}\psi_{1})$, equivalently
$\mathcal{H}_{0}(A^{+}_{1}\psi_{1})=E_{1}(A^{+}_{1}\psi_{1}).$ (7)
If $A^{+}_{1}\psi_{1}\neq 0$ then $E_{1}$ is also an eigenvalue of $\mathcal{H}_{0}$. The norm $\lVert A^{+}_{1}\psi_{1}\rVert=\sqrt{\braket{A^{+}_{1}\psi_{1}|A^{+}_{1}\psi_{1}}}$ is just $\sqrt{E_{1}-\mathscr{E}}$ and hence the normalized eigenfunction is
$\underset{\scriptsize\text{(normalized)}}{\psi_{0}}=\frac{A^{+}_{1}\psi_{1}}{\lVert A^{+}_{1}\psi_{1}\rVert}=\frac{1}{\sqrt{E_{1}-\mathscr{E}}}(A^{+}_{1}\psi_{1}).$ (8)
This is typically the situation unless $E_{1}$ is the ground state of $\mathcal{H}_{1}$, and this ensures the potentials have otherwise identical spectra. The interesting condition occurs when $A^{+}_{1}\psi_{1}=0$ which implies that $A^{-}_{1}A^{+}_{1}\psi_{1}=0$ and $\mathcal{H}_{1}\psi_{1}=\mathscr{E}\psi_{1}$. Then the eigenvalue $E_{1}=\mathscr{E}$ but $\psi_{0}$ is unnormalizeable since the norm in the denominator of (8) is zero. We then have a situation where the new potential $V_{1}(x)$ associated with Hamiltonian $\mathcal{H}_{1}$ is expected to have a ground state at energy $E_{1}<E_{0}$ which $V_{0}(x)$ lacks. As for $U_{1}(x)$, by combining (3), (5), and (6), it can be shown that it must satisfy a Ricatti equation,
@@
where prime denotes derivative with respect to $x$. The new potential is
$V_{1}(x)=V_{0}(x)-U^{\prime}_{1}(x)=U_{1}^{2}(x)-V_{0}(x)+2\mathscr{E}\,$, (10)
where Eq. (9) has been used. A general solution of Eq. (9) is given in [8] and [13] and has the form
$U_{1}(x)$ $=\frac{d}{dx}\left[\ln\psi_{0}(x)\right]\,.$ (11)
The unnormalized wavefunction $\psi_{0}(x)=A^{+}_{1}\psi_{1}$ pertains to the original function $V_{0}(x)$ and can be found by solving the Schrödinger equation (7) that we rewrite as
@@
It is convenient to split $\psi_{0}(x)$ into even and odd parts, $\psi_{0}(x)=\psi_{\mathrm{even}}(x)+\alpha\psi_{\mathrm{odd}}(x)$, with the appropriate initial conditions $\psi^{\prime}_{\mathrm{even}}(0)=0,\psi_{\mathrm{even}}(0)=1$ and $\psi_{\mathrm{odd}}(0)=0,\psi^{\prime}_{\mathrm{odd}}(0)=1$. Since the new energy level $\mathscr{E}$ lies below the ground state of $V_{0}(x)$, the requirement that $\mathscr{E}<E_{0}$ will ensure that $\psi_{\mathrm{even}}(x)$ be free of zeros. If $\psi_{\mathrm{odd}}(x)$ is included in the solutions of $\psi_{0}(x)$, care must be taken that $\alpha$ is not too large, otherwise the zero always present in $\psi_{\mathrm{odd}}(x)$ would show up in $\psi_{0}(x)$.
For a practical calculation of the new potential $V_{1}(x)$, it is more convenient to work with $U_{1}(x)$ rather than $\psi_{0}(x)$ because one avoids the astronomically large values that the unnormalizable $\psi_{0}(x)$ contains. To solve Eq. (9), we also need to employ the initial condition $U_{1}(0)=\alpha$ that follows from the initial conditions for $\psi_{0}$; the choice $\alpha=0$ yields a symmetric potential $V_{1}(x)$ and in the following we will restrict ourselves to this case.
@@
  • Starting potential $V_{0}(x)$
  • An array of desired new levels $E$ of size $m$ in descending order
Result: New potential $V_{m}(x)$
for $i\leftarrow 1$ to $m$ do
@@
 $U_{i}^{\prime}+U_{i}^{2}+2[E_{i}-V_{i-1}(x)]=0$,
   where $U_{i}(0)=0$.
   Create potential $V_{i}(x)$:
 $V_{i}(x)\leftarrow U_{i}^{2}(x)-V_{i-1}(x)+2E_{i}\,.$
return $V_{m}(x)$;
Output the final potential containing additional $m$ energy levels.
Algorithm 1 Procedure for Creating Potential with $m$ New Energy Levels
@@
Equation (10) provides the desired potential $V_{1}$ with the added level $\mathscr{E}$ and it is suitable for numerical calculations. The entire procedure can be iterated to add arbitrarily many new energy levels, provided that each new level be below all of the previously added levels; refer to Algorithm 1 for a complete description. We have tested the algorithm to construct several potentials (see below), and it can easily be used to add hundreds of levels.
@@
The intertwining method works, by its nature, in the “top to bottom” direction: we start from a given potential and add new energy levels below its existing levels. However, it might be more natural to work “bottom to top”, starting with the ground state and then adding sequentially higher energy levels. Fortunately, the method is suitable for this purpose too: We first list as many levels as desired that follow a required “bottom to top” pattern in our new potential. We then choose any suitable initial potential, typically the harmonic oscillator or a constant potential; we then invert our list of desired levels and add them in reverse order. Numerical simulations reveal that when many levels are added, then the choice of the initial potential becomes unimportant and does not influence the resulting potential on the relevant energy interval.
IV Constructing potentials with perfect revivals
In the following we introduce several examples of energy level structure that yield perfect quantum state revivals, and construct the corresponding potentials. We also demostrate the revivals using the autocorrelation function and quantum carpets.
IV.1 Biperiodic oscillator
Let’s start from the harmonic oscillator potential with energy levels $E_{n}=n+\frac{1}{2}$ for $n\geq 0$ and sequentially add an additional $N$ levels $E=-1,-3,-5,\ldots,-(2N-1)$. Employing the semiclassical relation between separation of quantum energy levels and classical period of oscillations [1], we find that this would correspond to a classical biperiodic oscillator that has frequency $\omega_{1}=1$ for $E>0$ and $\omega_{2}=2$ for $E<0$; such a potential was described for instance in Ref. [14]. The resulting quantum potentials and energy levels are shown in Fig. 1 for $N$ ranging from 1 to 25; the levels were numerically calculated from the constructed potentials and they match the desired levels with a high precision. The quantum potentials closely resemble their classical counterpart (see Fig. 1(a) in Ref. [14]). We can conveniently represent the time evolution and revival of the wavepacket by the quantum carpet [1], which is analogous to the Talbot carpet in optics [15]. Fig. 2 shows examples of quantum carpets that represent the time evolution of wavepackets for the biperiodic potential with $N=100$ added levels, where the bi-periodicity is apparent. Fig. 3(a) shows the magnitude of the autocorrelation function $|A(t)|$ and confirms perfect wavepacket revivals.

### v2

@@ Abstract
It is known that there exist a limited number of analytic potentials with the unusual property that any bound quantum state therein will be periodic in time. This is known as a perfect quantum state revival. Examples of such potentials are the infinite well, quantum harmonic oscillator and the Pöschl-Teller potentials; here, we present a general method of designing such potentials. A key requirement is that their energy eigenvalues have integer spacings (up to a prefactor). We first analyze the required conditions which permit quantum state revivals for potentials in general, and then we use techniques of iterated Hamiltonian intertwining to construct potentials exhibiting perfect quantum revivals. Our method can readily be extended to multiple dimensions.
@@ III Constructing potentials with desired spectra
In the following we present an adaptation of the standard intertwining procedure [10, 11, 8, 12]. The technique begins with a known potential $V_{0}(x)$ and associated Hamiltonian,
@@
A new Hamiltonian $\mathcal{H}_{1}$ corresponding to a potential $V_{1}(x)$ shall be constructed such that it has all the energy levels that $\mathcal{H}_{0}$ has, plus an additional level that shall lie below all those of $\mathcal{H}_{0}$. The total number of bound states of $\mathcal{H}_{1}(x)$ is thus one greater than in $\mathcal{H}_{0}$. If $V_{0}(x)$ is one of the potentials whose discrete levels obey the constraints of (2) and we then add a new energy level below its ground state at the right energy, a new potential retaining the quantum state revival property will result. It is then possible to iterate the intertwining procedure with $V_{1}(x)$ as the starting point to add another energy level below its ground state, and then continue to create as many new energy levels as desired.
To find $V_{1}(x)$, we first factorize the Hamiltonian $\mathcal{H}_{0}$ as
$\mathcal{H}_{0}=A^{+}A^{-}+\mathscr{E}\,$$\qquad$ (4)
where $A^{\pm}$ are mutually Hermitian-conjugated operators in the form
$A^{\pm}=\frac{1}{\sqrt{2}}\left[\pm\frac{d}{dx}+U_{1}(x)\right]$ (5)
and $U_{1}(x)$ is a function to be determined. The factorization (4) is possible if the value of $\mathscr{E}$ does not exceed the ground state energy $E_{0}^{(0)}$ of the Hamiltonian $\mathcal{H}_{0}$ because the operator $A^{+}A^{-}=\mathcal{H}_{0}$-$\mathscr{E}$ must be positive semidefinite. In the following we will assume that $\mathscr{E}<E_{0}^{(0)}$, and $\mathscr{E}$ will play the role of the new energy level to be added. Upon substituting Eqs. (5) and  (3) into Eq. (4) while taking care that the operators are handled correctly, we get an equation for $U_{1}(x)$
@@
where prime denotes derivative with respect to $x$. Equation (6) is known as Riccati equation [13, 14].
The new Hamiltonian with the added level $\mathscr{E}$ will then be
$\mathcal{H}_{1}$ $=A^{-}A^{+}+\mathscr{E}\,.$ (7)
To show that $\mathcal{H}_{1}$ has the required properties, we have to show that (i) all the energy levels of $\mathcal{H}_{0}$ are also energy levels of $\mathcal{H}_{1}$, and (ii) $\mathcal{H}_{1}$ has an additional level $\mathscr{E}$ that $\mathcal{H}_{0}$ does not have.
To show (i), we denote by $\psi$ a normalized eigenstate of the Hamiltonian $\mathcal{H}_{0}$ with energy $E$. Then Eq. (4) yields $A^{+}A^{-}\psi=(E-\mathscr{E})\psi$. Multiplying from the left by $A^{-}$ results in $A^{-}A^{+}(A^{-}\psi)=(E-\mathscr{E})(A^{-}\psi)$, or, equivalently,
$\mathcal{H}_{1}(A^{-}\psi)=E(A^{-}\psi)\,$, (8)
where Eq. (7) was used. The square of the norm of the function $A^{-}\psi$ can readily be calculated as $\|A^{-}\psi\|^{2}=\langle A^{-}\psi|A^{-}\psi\rangle=\langle\psi|A^{+}A^{-}|\psi\rangle=E-\mathscr{E}>0$, which can be used to normalize the function $A^{-}\psi$ as $\psi_{1}\equiv(A^{-}\psi)/\sqrt{E-\mathscr{E}}$. From Eq. (8) we then see that $\mathcal{H}_{1}\psi_{1}=E\psi_{1}$, i.e., the new Hamiltonian $\mathcal{H}_{1}$ has a normalized wavefunction $\psi_{1}$ with energy $E$, so it shares the eigenvalue $E$ with the old Hamiltonian $\mathcal{H}_{1}$.
To show (ii), we take advantage of the fact that Eq. (6) has the solution of the form [8, 14]
$U_{1}(x)$ $=\frac{d}{dx}\left[\ln\psi_{0}(x)\right]=\frac{\psi^{\prime}_{0}(x)}{\psi_{0}(x)}\,$, (9)
where the function $\psi_{0}(x)$ satisfies
@@
This can easily be verified by substituting Eq. (9) into Eq. (6) and employing Eq. (10). At the same time, Eq. (10) is nothing else than the stationary Schrödinger equation $\mathcal{H}_{0}\psi_{0}=\mathscr{E}\psi_{0}$. The function $\psi_{0}$ cannot be normalizable, however, since $\mathscr{E}$ does not belong to the spectrum of $\mathcal{H}_{0}$; as we have assumed, $\mathscr{E}$ lies below the ground state energy of $\mathcal{H}_{0}$. Still, $\psi_{0}$ turns out to be very useful because it can be used for constructing the desired normalizable eigenstate of $\mathcal{H}_{1}$ with eigenvalue $\mathscr{E}$.
To do that, we realize that the function $\psi_{0}$ must be nodeless, otherwise Eq. (9) would lead to a singular function $U_{1}(x)$. A general solution of Eq. (10) contains two arbitrary constants that can be chosen as $\alpha=\psi^{\prime}_{0}(0)$ and $\beta=\psi_{0}(0)$. Equation (9) further reveals that multiplying $\psi_{0}$ by a constant does not change $U_{1}$, so without loss of generality we can assume that $\beta=1$.
It is convenient to split $\psi_{0}(x)$ into even and odd parts, $\psi_{0}(x)=\psi_{\mathrm{even}}(x)+\alpha\psi_{\mathrm{odd}}(x)$, with the initial conditions $\psi^{\prime}_{\mathrm{even}}(0)=0,\psi_{\mathrm{even}}(0)=1$ and $\psi_{\mathrm{odd}}(0)=0,\psi^{\prime}_{\mathrm{odd}}(0)=1$. The fact that $\mathscr{E}$ lies below the ground state of $\mathcal{H}_{0}(x)$ ensures that $\psi_{\mathrm{even}}(x)$ be free of zeros. If $\psi_{\mathrm{odd}}(x)$ is included in the solutions of $\psi_{0}(x)$, care must be taken that $|\alpha|$ is not too large, otherwise the zero present in $\psi_{\mathrm{odd}}(x)$ would show up in $\psi_{0}(x)$. On the other hand, if $\alpha$ is chosen appropriately, then $\psi_{0}(x)$ is positive for all $x$, which then yields a non-singular, real function $U_{1}(x)$ via Eq. (9). Moreover, due to the fact that $\mathscr{E}$ lies below the ground state of $\mathcal{H}_{0}(x)$, the function $\psi_{0}(x)$ will have an exponentially growing behavior for $x\to\pm\infty$; the only exception would be choosing a particular positive value of $\alpha$ such that $\psi_{0}(x)$ would converge to zero for $x\to$-$\infty$ (but still diverge for $x\to\infty$), or choosing a particular negative value of $\alpha$ such that $\psi_{0}(x)$ would converge to zero for $x\to\infty$ (but still diverge for $x\to$-$\infty$). Excluding such values from the possible choices of $\alpha$, we arrive at a positive real function $\psi_{0}(x)$ that has an exponentially growing behavior for $x\to\pm\infty$. Now consider a function $\psi_{01}(x)=1/\psi_{0}(x)$; it is normalizable due to the exponentially growing behavior of $\psi_{0}(x)$. A simple calculation further reveals that
$\sqrt{2}\,A^{+}\psi_{01}=\psi^{\prime}_{01}+U_{1}\psi_{01}=-\frac{\psi^{\prime}_{0}}{\psi^{2}_{0}}+\frac{\psi^{\prime}_{0}}{\psi_{0}}\frac{1}{\psi_{0}}=0\,.$ (11)
From this it follows that $A^{-}A^{+}\psi_{01}=(\mathcal{H}_{1}-\mathscr{E})\psi_{01}=0$; in other words, the Hamiltonian $\mathcal{H}_{1}$ has a normalizable wavefunction with the eigenvalue $\mathscr{E}$, which completes the part (ii).
The last step to be done is expressing the potential $V_{1}(x)$ in the new Hamiltonian $\mathcal{H}_{1}$. This can easily be performed by combining Eqs. (7) and (5), which yields
$\mathcal{H}_{1}$ $=\frac{1}{2}\left[-\frac{d}{dx}+U_{1}(x)\right]\left[\frac{d}{dx}+U_{1}(x)\right]+\mathscr{E}$
$=-\frac{1}{2}\frac{d^{2}}{dx^{2}}+V_{1}(x)\,$, (12)
where the new potential is
$V_{1}(x)=\frac{1}{2}\,[U_{1}^{2}(x)-U^{\prime}_{1}(x)]+\mathscr{E}=U_{1}^{2}(x)-V_{0}(x)+2\mathscr{E}$ (13)
and Eq. (6) has been used. Alternatively, $V_{1}(x)$ could be expressed in terms of the function $\psi_{0}$, which leads to
$V_{1}(x)=V_{0}(x)-\frac{d^{2}}{dx^{2}}\,\ln\psi_{0}\,.$ (14)
This is the form derived in Ref. [8]. It reveals that due to the properties of the function $\psi_{0}(x)$ discussed above, $V_{1}(x)$ is a real and well behaved function.
For a practical calculation of the new potential $V_{1}(x)$, in particular a numerical calculation, it is more convenient to use Eq. (13) rather than Eq. (14) because one then avoids the astronomically large values that the unnormalizable $\psi_{0}(x)$ contains. The $U_{1}(x)$ is found by solving Eq. (6), for which an initial condition is also needed; it follows from the initial conditions for $\psi_{0}(x)$ and has the form $U_{1}(0)=\alpha$. Provided that the original potential is symmetric, i.e., $V_{0}(-x)=V_{0}(x)$, the choice $\alpha=0$ yields an antisymmetric (odd) function $U_{1}(x)$ and consequently a symmetric potential $V_{1}(x)$; in the following we will restrict ourselves to this case.
@@
  • Starting symmetric potential $V_{0}(x)$
  • A decreasing sequence of desired new levels $\mathscr{E}_{1},\dots,\mathscr{E}_{n}$
Result: New potential $V_{n}(x)$
Solve the equation for $U_{1}(x)$:
$U^{\prime}_{1}(x)+U_{1}^{2}(x)=2[V_{0}(x)-\mathscr{E}_{1}]$
with $U_{1}(0)=0$.
for $i\leftarrow 2$ to $n$ do
@@
 $U^{\prime}_{i}+U^{2}_{i}=-U^{\prime}_{i-1}+U^{2}_{i-1}+2(\mathscr{E}_{i-1}-\mathscr{E}_{i})$
   with $U_{i}(0)=0$.
return $V_{n}=(U_{n}^{2}-U_{n}^{\prime})/2+\mathscr{E}_{n}$;
Output the final potential containing the additional $n$ energy levels.
Algorithm 1 Procedure for Creating Potential with $n$ New Energy Levels
@@
Equation (13) provides the desired new potential $V_{1}(x)$ with the added level $\mathscr{E}$ and it is suitable for numerical calculations. The entire procedure can be iterated to add arbitrarily many new energy levels, provided that each new level be below all of the previously added levels. This way, if $n$ new levels $\mathscr{E}_{1},\dots,\mathscr{E}_{n}$ are being added consecutively, one obtains a sequence of functions $U_{1},\dots,U_{n}$ and potentials $V_{1},\dots,V_{n}$. However, for a practical calculation when only the last potential $V_{n}$ is of interest, it is more convenient to work directly with the functions $U_{i}$ and avoid the functions $V_{i}$ in the intermediate steps. For this purpose, one can derive a recurrence equation for the functions $U_{i}$ in the form
$U^{\prime}_{i}(x)+U^{2}_{i}(x)=-U^{\prime}_{i-1}(x)+U^{2}_{i-1}(x)+2(\mathscr{E}_{i-1}-\mathscr{E}_{i})$ (15)
for $i=2,3,\dots,n$. For calculating $U_{1}$, Eq. (6) is used with $\mathscr{E}$ replaced by $\mathscr{E}_{1}$, and the final potential is obtained as
$V_{n}(x)=\frac{1}{2}[U_{n}^{2}(x)-U_{n}^{\prime}(x)]+\mathscr{E}_{n}$ (16)
in analogy with Eq. (13). Algorithm 1 summarizes the whole procedure. We have tested the algorithm to construct potentials with various spectral properties; a number of examples are shown in the next section. While it is known that this method can be used to add energy levels to existing potentials (e.g., a single level insertion is illustrated in Ref.  [8]), we have found that hundreds of levels can be added easily when working with $U_{i}$.
Aside from this caveat, however, the intertwining method itself is well known, with a formalism similar to that of supersymmetric quantum mechanics [15] which emerged from Witten’s pioneering work [16] as a simplified framework to study symmetry breaking in quantum field theory, later revealing deep connections with solvable quantum potentials. In this formalism, $A^{-}$ and $A^{+}$ are known as supercharge operators and $U_{1}(x)$ is the superpotential. The partner Hamiltonians $H_{0}$ and $H_{1}$ are said to be intertwined by these operators. Historically, these constructions trace back to Darboux’s transformation [17], which described how one linear differential equation could generate another with shifted spectral properties; this was later systematized in quantum mechanics [18, 19]. Mielnik [20] generalized the factorization to yield a continuous set of isospectral deformations, and showed this in depth with the example of the quantum harmonic oscillator. Further developments by Andrianov et al. [21] and Bagrov and Samsonov [22] established the correspondence between Darboux transformations, intertwining operators, and supersymmetric partner Hamiltonians. These works showed that successive Darboux transformations could generate new potentials with selectively added or removed bound states, which we employ precisely for such spectral engineering. This demonstrates the applicability of supersymmetry far beyond its original intentions. Indeed, it can be said that the very existence of potentials with identical spectra is due to hidden supersymmetry [21].
The intertwining method works, by its nature, in the “top to bottom” direction: we start from a given potential and add new energy levels below its existing levels. However, it might be more natural to work “bottom to top”, starting with the ground state and then adding sequentially higher energy levels. Fortunately, the method is suitable for this purpose too: We first list as many levels as desired that follow a required “bottom to top” pattern in our new potential. We then choose any suitable initial potential, typically the harmonic oscillator or a constant potential; we then invert our list of desired levels and add them in reverse order.
III.1 Influence of the choice of $V_{0}(x)$ on $V_{n}(x)$
It is natural to ask what is the influence of the choice of the original potential $V_{0}(x)$ on the resulting final potential $V_{n}(x)$ for given energy levels $\mathscr{E}_{1},\dots,\mathscr{E}_{n}$ to be added. While it turns out that the choice of $V_{0}$ influences $V_{n}$ in complex ways, interestingly, the value of $V_{n}$ at the origin can be calculated analytically. By combining Eq. (6), Eqs. (15) for $i=2,\dots,n$ and Eq. (16), and employing the fact that $U_{i}(0)=0$ for all $i=1,\dots,n$, we arrive at
$V_{n}(0)=(-1)^{n}\,V_{0}(0)+2\sum_{i=1}^{n}(-1)^{n-i}\,\mathscr{E}_{i}\,.$ (17)
This equation shows that if we keep the added levels $\mathscr{E}_{1},\dots,\mathscr{E}_{n}$ fixed and increase the value $V_{0}(0)$ [e.g., by shifting the whole potential $V_{0}(x)$], the value $V_{n}(0)$ increases by the same amount if $n$ is even while it decreases by the same amount if $n$ is odd. Consequently, choosing the original potential $V_{0}$ inappropriately can lead to strong oscillations of the potential $V_{n}$.
It is in principle possible to evaluate also the derivatives $V_{n}^{(k)}(0)$ analytically; for odd $k$, they are identically zero while for even $k$ they lead to increasingly complex expressions involving the added energy levels and the derivatives of the original potential $V_{0}$ at the origin.
IV Constructing potentials with perfect revivals
In the following we introduce several examples of energy level structure that yield perfect quantum state revivals, and construct the corresponding potentials. We also demonstrate the revivals using the autocorrelation function and quantum carpets.
IV.1 Biperiodic oscillator
@@
Let’s start from the harmonic oscillator potential with energy levels $E_{n}=n+\frac{1}{2}$ for $n\geq 0$ and sequentially add an additional $N$ levels $E=-1,-3,-5,\ldots,-(2N-1)$. Employing the semiclassical relation between separation of quantum energy levels and classical period of oscillations [1], we find that this would correspond to a classical biperiodic oscillator that has frequency $\omega_{1}=1$ for $E>0$ and $\omega_{2}=2$ for $E<0$; such a potential was described for instance in Ref. [23]. The resulting quantum potentials and energy levels are shown in Fig. 1 for $N$ ranging from 1 to 25; the levels were numerically calculated from the constructed potentials and they match the desired levels with a high precision. The quantum potentials closely resemble their classical counterpart (see Fig. 1(a) in Ref. [23]). We can conveniently represent the time evolution and revival of the wavepacket by the quantum carpet [1], which is analogous to the Talbot carpet in optics [24]. Such a carpet represents the time evolution of a wavepacket $\psi(x,t)$ as a 2D plot of the probability density $|\psi(x,t)|^{2}$, where the time variable is on the vertical axis and coordinate on the horizontal axis. Perfect wavepacket revivals manifest themselves by the periodicity of the carpet in the time direction.
Fig. 2 shows examples of quantum carpets that represent the time evolution of wavepackets for the biperiodic potential with $N=100$ added levels, where the bi-periodicity is apparent. Fig. 3(a) shows the magnitude of the autocorrelation function $|A(t)|$ and confirms perfect wavepacket revivals.